# WidthFormer: Toward Efficient Transformer-based BEV View Transformation

### v3

@@ 4.3 Ablation Studies and Discussions
Building WidthFormer step by step.
We first examine the effect of our proposed modules by gradually adding them to the model and reporting accuracy & speed results in Tab. 3. We start with the original LSS [23] baseline that achieves 29.5 mAP with 4.5 ms VT latency. We then replace LSS with a transformer layer, where BEV queries directly interact with the image features without 3D positional encoding (PE). However, the model fails to converge. We then add our RefPE encoding to the model, in which we compute PEs for every feature pixel without averaging them in the height dimension. The model achieves 31.0 mAP, which is better than the LSS baseline. This result validates the necessity of 3D PEs. However, querying the whole image features is inefficient (4.6 ms). By adopting width features as attention keys and values, the latency reduces to 1.3 ms. However, the accuracy falls to 28.3 mAP, implying that the feature compression is causing us to lose information. We mitigate this by adding the Refine Transformer to the model, which brings 1.0 mAP improvement with minor computation overhead (+0.2 ms). Finally, we add the complementary tasks that improve the final mAP to 30.7. These results show that the proposed Refine Transformer and complementary tasks can indeed mitigate the information loss problem.
Refine Transformer.
@@
We examine the design of our Refine Transformer and present the results in Tab. 4. The baseline model that has no refinement to the width features achieves 30.0 mAP and 1.3 ms VT latency. When adding the convolutional-based refinement in MatrixVT [42], the accuracy is improved to 30.2 mAP. However, when adding our proposed Refine Transformer, the mAP is greatly improved to 30.7. We notice that our Refine Transformer is slightly slower than the convolution-based refinement but that this is negligible compared to the overall latency. We also tried to combine Refine Transformer with convolution-based refinement but found it only brought a marginal improvement in accuracy. This experiment validates the effectiveness of our Refine Transformer.
Complementary Tasks.
@@
We study our proposed complementary tasks and report the results in Tab. 5. When training with complementary tasks is disabled, our model achieves 29.3 mAP. Then we add the 3D monocular detection task as in FCOS3D [33], the detection accuracy is improved to 29.7 mAP. Then we add the height estimation branch—where the model estimates an object’s height in the 2D image feature—the performance is further improved by 30.5 mAP. Finally, we modified the regression depth estimation loss into a categorical depth estimation loss, which gave us a final mAP to 30.7. This experiment validates the effectiveness of our training strategy.
Speed on Edge-computing chip.
@@
In Fig. 6, we report the speed-testing results comparing WidthFormer and other VT methods on the Horizon Journey-5 [24] edge-computing chip. We compare our method with three other competing methods: GKT [3], LSS [23] and MatrixVT [42]. We report the VT latency with different size settings using square BEV grids. Specifically, the five size settings are: S=, S=, S=, S= and S=. Note that the Journey-5 chip has special support for LSS, making it much faster than naive implementations. The results show that WidthFormer is nearly two times faster than LSS and several times faster than GKT. We also notice that MatrixVT has a similar latency to ours, which differs from the results in Fig. 4. After investigation, we found that MatrixVT’s onboard speed is severely limited by the two very large ray and ring matrices. On the other hand, as shown in other experiments, WidthFormer has a better accuracy (Tab. 2) than MatrixVT. Therefore, we argue that ours is a better choice for the actual autonomous driving applications.
@@
Robustness under camera perturbation. In this section, we study the robustness of our method and other VT approaches under 6DoF camera perturbation: unexpected rotation and translation w.r.t. , and axis in the camera coordinate system. Specifically, The perturbations are implemented by modifying the camera poses with zero-mean Gaussian noises, whose magnitude is controlled by the standard deviation of the added Gaussian. We independently study the types of perturbations. We show the different VT methods’ mAP curves w.r.t. different types of perturbation in Fig. 5. Interestingly, we find that the different VT methods are sensitive to different types of perturbations. For example, while the -axis translation causes performance decrease to IPM and BEVFormer, our method and Lift-splat based LSS and MatrixVT are completely insensitive to this perturbation because the height dimension is omitted in these approaches. In general, we observe that the Lift-splat based approaches are more robust than IPM and BEVFormer. The reason is that IPM and BEVFormer rely on projecting 3D space coordinates to 2D images, and a small amount of perturbation may cause a huge difference in the projected location, therefore harming performance. Particularly, our method is very robust to the translation perturbations with minor performance degrades compared to all competing approaches. Our method is also robust against and axis rotation. However, we notice that all VT methods, including ours, have similar strong sensitivities to the -axis rotation, with the mAPs rapidly decreasing to zero as the noise magnitudes rise. The reason is that the -axis rotation can severely damages the predefined multi-view arrangement [1].
@@
Scaling-up Experiments. In Tab. 6, we scale up our model and compare it with other state-of-the-art 3D detection approaches on the nuScenes val set. Specifically, for BEV-based detectors, we use ResNet-101 as the image encoder; we also scale up the input resolution and BEV channel dimensions to 5121408 and 128, respectively. The results show that our method achieves on-par performance with other state-of-the-art methods. For example, BEVFormer [18] achieves 37.5 mAP & 44.8 NDS and 41.6 mAP & 51.7 NDS in single-frame and multi-frame settings, while ours achieves 37.9 mAP & 44.8 NDS and 40.1 mAP & 52.7 NDS. In the multi-modality setting, the recently proposed BEVDistill [6] achieves 41.7 mAP & 52.4 NDS, while ours achieves 40.9 mAP & 53.3 NDS. We also compare our StreamPETR+RefPE with other high-performing methods using the V2-99 backbone, which also validates our method’s superiority.
5 Conclusion
@@
This work introduces WidthFormer, a new transformer-based BEV transformation method. With the help of a new 3D positional encoding RefPE, which can also be used to boost a sparse 3D detector’s performance, WidthFormer use a single layer of transformer decoder to compute the BEV representations from image features, which are compressed in the vertical direction. Compared with previous BEV transformation approaches, our model is more efficient and easy to deploy to edge-computing devices. We also show that WidthFormer holds a good robustness against camera perturbations. We hope WidthFormer can serve as a simple baseline model for future research and real-world autonomous driving applications.

### v5

@@ IV-C Ablation Studies and Discussions
Refine Transformer. We examine the design of our Refine Transformer and present the results in Tab. IV. The baseline model that has no refinement to the width features achieves 30.0 mAP and 1.3 ms VT latency. When adding the convolutional-based refinement in MatrixVT [36], the accuracy is improved to 30.2 mAP. However, when adding our proposed Refine Transformer, the mAP is greatly improved to 30.7. We notice that our Refine Transformer is slightly slower than the convolution-based refinement but that this is negligible compared to the overall latency. We also tried to combine Refine Transformer with convolution-based refinement but found it only brought a marginal improvement in accuracy. This experiment validates the effectiveness of our Refine Transformer.
@@
Complementary Tasks. We study our proposed complementary tasks and report the results in Tab. V. When training with complementary tasks is disabled, our model achieves 29.3 mAP. Then we add the 3D monocular detection task as in FCOS3D [29], the detection accuracy is improved to 29.7 mAP. Then we add the height estimation branch—where the model estimates an object’s height in the 2D image feature—the performance is further improved by 30.5 mAP. Finally, we modified the regression depth estimation loss into a categorical depth estimation loss, which gave us a final mAP to 30.7. This experiment validates the effectiveness of our training strategy.
@@
Speed on Edge Computing Device. In Fig. 5, we report the speed-testing results comparing WidthFormer and other VT methods on the Horizon Journey-5 [22]. We compare our method with three other competing methods: GKT [3], LSS [21] and MatrixVT [36]. We report the VT latency with different size settings using square BEV grids. Specifically, the five size settings are: S1=, S2=, S3=, S2= and S5=. Note that the Journey-5 has special support for LSS, making it much faster than naive implementations. The results show that WidthFormer is nearly two times faster than LSS and several times faster than GKT. We also notice that MatrixVT has a similar latency to ours, which differs from the results in Fig. 4. After investigation, we found that MatrixVT’s onboard speed is severely limited by the two very large ray and ring matrices. On the other hand, as shown in other experiments, WidthFormer has a better accuracy (Tab. II) than MatrixVT. Therefore, we argue that ours is a better choice for the actual autonomous driving applications.
@@
Robustness under camera perturbation. In this section, we study the robustness of our method and other VT approaches under 6DoF camera perturbation: unexpected rotation and translation w.r.t. , and axis in the camera coordinate system. Specifically, the perturbations are implemented by modifying the camera extrinsics with zero-mean Gaussian noises, whose magnitude is controlled by the standard deviation of the added Gaussian. Specifically, for the -th camera, the noised rotation matrix is computed by , and the noised translation vector is computed by , where denotes matrix multiplication; and are the noisy rotation matrix and translation vector. We then use or to compute Eq. 1. We show the different VT methods’ mAP curves w.r.t. different types of perturbation in Fig. 6. Interestingly, we find that the different VT methods are sensitive to different types of perturbations. For example, while the -axis translation causes performance decrease to IPM and BEVFormer, our method and Lift-splat based LSS and MatrixVT are completely insensitive to this perturbation because the height dimension is omitted in these approaches. In general, we observe that the Lift-splat based approaches are more robust than IPM and BEVFormer. The reason is that IPM and BEVFormer rely on projecting 3D space coordinates to 2D images, and a small amount of perturbation may cause a huge difference in the projected location, therefore harming performance. Particularly, our method is very robust to the translation perturbations with minor performance degrades compared to all competing approaches. Our method is also robust against and axis rotation. However, we notice that all VT methods, including ours, have similar strong sensitivities to the -axis rotation, with the mAPs rapidly decreasing to zero as the noise magnitudes rise. The reason is that the -axis rotation can severely damages the predefined multi-view arrangement [1].
@@
Scaling-up Experiments. In Tab. VI, we scale up our model and compare it with other state-of-the-art 3D detection approaches on the nuScenes val set. Specifically, for BEV-based detectors, we use ResNet-101 as the image encoder; we also scale up the input resolution and BEV channel dimensions to 5121408 and 128, respectively. The results show that our method achieves on-par performance with other state-of-the-art methods. For example, BEVFormer [16] achieves 37.5 mAP & 44.8 NDS and 41.6 mAP & 51.7 NDS in single-frame and multi-frame settings, while ours achieves 37.9 mAP & 44.8 NDS and 40.1 mAP & 52.7 NDS. Additionally, the superior performance of StreamPETR+RefPE further validates the effectiveness of our design.
V Conclusion
@@
We introduce WidthFormer, a new transformer-based BEV transformation method. With the help of a new 3D positional encoding RefPE, which can also be used to boost a sparse 3D detector’s performance, WidthFormer use a single layer of transformer decoder to compute the BEV representations from the vertically compressed visual features. Compared with previous BEV transformation approaches, our model is more efficient and easy to deploy to edge-computing devices. We also show that WidthFormer holds a good robustness against camera perturbations.